# Lower bounds for Seshadri constants on blow Ups of $\mathbb{P}^{2}$

### v2

@@ 3. Seshadri constants on Blow ups of $r\leq 8$ distinct points
[11, Proposition 2.5] Let $X_{r}$ be the blow up of $r$ distinct points in $\mathbb{P}^{2}$ and let $L$ denote an ample line bundle on $X_{r}$. Assume $r\leq 8$, and for $x\in X_{r}$, let $\widetilde{X}_{r}(x)$ denote the blow up of $X_{r}$ at $x$. Then
@@
where $\Gamma$ is the set of all $(-1)$-curves, $(-2)$-curves, fixed components of $\big{|}-K_{\widetilde{X}_{r}(x)}\big{|}$ or reduced irreducible curves in the linear system $\big{|}-K_{\widetilde{X}_{r}(x)}\big{|}$ on $\widetilde{X}_{r}(x)$.
@@
If $C$ is a Seshadri curve in $X_{r}$ for a line bundle at a point $x\in X_{r}$, then $\widetilde{C}^{2}<0$, where $\widetilde{C}$ is the strict transform of $C$ on $\widetilde{X}_{r}(x)$. Note that by [17, Proposition 4.1 (ii)], for $r\leq 8$, $\widetilde{C}$ is either $(-1)$-curve, $(-2)$-curve, or a fixed component of $\big{|}-K_{\widetilde{X}_{r}(x)}\big{|}$.
@@
Let $X$ be a smooth anticanonical surface (i.e., $-K_{X}$ is an effective divisor). If $(-K_{X})^{2}\geq 2$ then for every $x\in X$,
$\varepsilon(X,L,x)\geq 1$
for any ample line bundle $L$ on $X$.
@@
As an application of the Hodge index theorem we have,
@@
So $-K_{X}\cdot L\geq\sqrt{2}$. Also note that $-K_{X}\cdot L$ is a positive integer. Thus we have $-K_{X}.L\geq 2$. Therefore by [13, Theorem III (a)], $L$ is base point free. So $\varepsilon(X_{r},L,x)\geq 1$ for all $x\in X_{r}$. ∎
@@
For $r\leq 7$, let $L$ be an ample line bundle on $X_{r}$. Then
$\varepsilon(X_{r},L,x)\geq 1$
for all $x\in X_{r}$.
Next, we consider the $r=8$ case. We first prove the following lemma.
@@
Let $X_{8}$ be a blow up of $\mathbb{P}^{2}$ at 8 distinct points. If $L$ is an ample line bundle with $-K_{X_{8}}\cdot L=1$ then $L=-K_{X_{8}}$.
@@
Let $L=eH-n_{1}E_{1}-\dots-n_{8}E_{8}$ be an ample line bundle with $-K_{X_{8}}\cdot L=1$, i.e., $13e-\displaystyle\sum_{i=1}^{8}n_{i}=$ 1. If $L^{2}\geq 2$, the Hodge index theorem gives
@@
Let $X_{8}$ be a blow up of $\mathbb{P}^{2}$ at 8 distinct points $p_{1},p_{2},\dots,p_{8}$. Let $L$ be an ample line bundle on $X_{8}$. Suppose that $\varepsilon(X_{8},L,x)<1$ for some $x\in X_{8}$. Then the following must hold:
@@
Suppose that $\varepsilon(X_{r},L,x)<1$ for an ample line bundle $L$ at some $x\in X_{r}$. By Proposition 3.1, the Seshadri constant is achieved by a curve $C$ on $X_{r}$ whose strict transform on $\widetilde{X}_{r}(x)$ is either a $(-1)$-curve or a $(-2)$-curve or a fixed component of $|-K_{\widetilde{X}_{r}(x)}|$ or a curve in the linear system $|-K_{\widetilde{X}_{r}(x)}|$. Let $C=dH-m_{1}E_{1}-\dots-m_{r}E_{r}$ and $t=\text{mult}_{x}C>1$, i.e,
@@
First, we show that the strict transform of $C$ on $\widetilde{X}_{r}(x)$ is not a $(-1)$ or a $(-2)$-curve. If $C$ is such a curve, we have $C^{2}=t^{2}-i$, where $i=1,2$. Since
@@
So $C$ is a singular curve whose strict transform is either a fixed component of $|-K_{\widetilde{X}_{r}(x)}|$ or a curve in the linear system $|-K_{\widetilde{X}_{r}(x)}|$. Since $C$ is a singular curve, we must have $d=3$. Hence $\widetilde{C}\in|-K_{\widetilde{X}_{r}(x)}|$. So $p_{1},\dots,p_{8}$ belong to the image of $C$ in $\mathbb{P}^{2}$. As $C$ is irreducible, it is clear that $\text{mult}_{x}C=2$. Since $\varepsilon(X_{8},L,x)<1$, the only possibility is
@@
A part of Theorem 3.6 has an alternate proof using Theorem 2.1. Note that there exists a cubic passing through any distinct 8 points in $\mathbb{P}^{2}$. Theorem 2.1 says that $\frac{1}{2}$ is a Seshadri constant of $L$ only if $L\cdot-K_{X_{8}}=1$ and Lemma 3.5 enables us to conclude that it can happen only if $L=-K_{X_{8}}$.
@@
By Theorem 3.6, the least Seshadri constant that can be achieved on $X_{8}$ is $\frac{1}{2}$. We now give an example in which $\frac{1}{2}$ is achieved at some point on a del Pezzo surface. Consider the smooth rational surface obtained by blowing up 8 smooth points $p_{1},p_{2},\dots p_{8}$ on a singular irreducible cubic in $\mathbb{P}^{2}$ such that no three lie on a line, no six lie on a conic, and if a singular cubic passes through all the points then $p_{i}$ are smooth points of that cubic for $1\leq i\leq 8$. Then $-K_{X_{8}}$ is ample by [4, Théorème 1]. Let $x$ denote the point in $X_{8}$ corresponding to the singular point on the cubic. Let $C$ denote the strict transform of the singular cubic containing the 8 points. From Theorem 3.6 it is clear that $\varepsilon(X_{8},-K_{X_{8}},x)\geq\frac{1}{2}$. Note that $\text{mult}_{x}C=2$ and $-K_{X_{8}}\cdot C=1$. These give us:
@@
It is known that the number of $(-1)$ and $(-2)$-curves in $\widetilde{X}_{r}(x)$ is finite for $r\leq 7$ and we can see that curves whose strict transforms are $(-1)$ or $(-2)$-curves are smooth in $\widetilde{X}_{r}(x)$ for $r\leq 5$. The curves in $X_{r}$ whose strict transform is a fixed component of $|-K_{\widetilde{X}_{r}(x)}|$ and is not an element of $|-K_{\widetilde{X}_{r}(x)}|$ are also smooth. It is also clear that a curve $C$ in $X_{r}$ such that $\widetilde{C}$ is an element of $|-K_{\widetilde{X}_{r}(x)}|$ has non-negative self intersection if $r\leq 5$. Note the fact that if $C$ is a Seshadri curve, its strict transform $\widetilde{C}$ has negative self intersection in $\widetilde{X}_{r}(x)$. This shows that all possible Seshadri curves are smooth when $r\leq 5$ which in fact implies that the Seshadri constant of every ample line bundle at every point of $X_{r}$ is an integer when $r\leq 5$. ∎
@@
It is shown in [2, Théorème 1.3] that when $r=6$, a fractional Seshadri constant is achieved. More precisely, $\varepsilon(X_{6},-K_{X_{6}},x)=\frac{3}{2}$ for some $x\in X_{6}$.
@@ 4. General blow ups of $\mathbb{P}^{2}$
We proceed with a similar setting as in the above section. Furthermore throughout this section we assume that $p_{1},p_{2},\dots,p_{r}$ are very general points in $\mathbb{P}^{2}$, where $r\geq 1$. We will give some lower bounds on Seshadri constants assuming the Strong SHGH Conjecture (Conjecture 1.3) is true.
@@
For any point $x\in X_{r}$, let $\widetilde{X}_{r}(x)$ be the blow up of $X_{r}$ at x. Assume that Conjecture 1.3 is true. Let $C$ be a reduced, irreducible curve on $X_{r}$ and let $\widetilde{C}$ denote its strict transform on $\widetilde{X}_{r}(x)$. If $\widetilde{C}^{2}<0$ then $\widetilde{C}$ is one of the following types.
@@
Let $C=dH-m_{1}E_{1}-\dots-m_{r}E_{r}$ and $t=\text{mult}_{x}C$. So $\widetilde{C}=dH-m_{1}E_{1}-\dots-m_{r}E_{r}-tE_{x}$, where $E_{x}$ is the exceptional divisor in $\widetilde{X}_{r}(x)$. If $d=0$, $\widetilde{C}=E_{i}$ or $\widetilde{C}=E_{i}-E_{x}$. Both are of the required type.
@@
(${d+2\choose 2}-\displaystyle\sum_{i=1}^{r}{m_{i}+1\choose 2}>{t+1\choose 2}-2.$
@@
which is the same as
$\frac{\widetilde{C}^{2}-K_{\widetilde{X}_{r}(x)}.\widetilde{C}}{2}+1>-2.$ (4.1)
The Reimann-Roch theorem gives,
@@
As $\widetilde{C}$ is an effective divisor in a smooth projective surface $\widetilde{X}_{r}(x)$, we have $h^{2}(\widetilde{X}_{r}(x),\widetilde{C})=0$. Since $\widetilde{C}^{2}<0$, $h^{0}(\widetilde{X}_{r}(x),\widetilde{C})=1$. Then by (4.1),
$-h^{1}(\widetilde{X}_{r}(x),\widetilde{C})=\frac{\widetilde{C}^{2}-K_{\widetilde{X}_{r}(x)}\cdot\widetilde{C}}{2}>-3.$
@@
    Then $h^{1}(\widetilde{X}_{r}(x),\widetilde{C})=\frac{K_{\widetilde{X}_{r}(x)}\cdot\widetilde{C}-\widetilde{C}^{2}}{2}=0$. So $K_{\widetilde{X}_{r}(x)}\cdot\widetilde{C}=\widetilde{C}^{2}$.
    As $\widetilde{C}$ is reduced and irreducible, we can apply the adjunction formula:
@@
    Similarly as in above cases, we conclude that types (3) or (5) or (6) can occur
The following is the main result of this section.
@@
Assume that Conjecture 1.3 is true. Let $\pi:X_{r}\to\mathbb{P}^{2}$ be the blow up of $\mathbb{P}^{2}$ at very general points $p_{1},p_{2},\dots,p_{r}$. Then for every ample line bundle L on $X_{r}$ and for all $x\in X_{r}$,
$\varepsilon(X_{r},L,x)\geq\frac{1}{2}.$
Further if $\varepsilon(X_{r},L,X)<1$ then $\varepsilon(X_{r},L,x)=\frac{1}{2}$.
@@
Suppose that $\varepsilon(X_{r},L,x)<1$ for some ample line bundle $L$ and some $x\in X_{r}$. Then the Seshadri constant is achieved by an irreducible and reduced curve $C\subseteq X_{r}$, i.e.,
@@
Let $L=eH-n_{1}E_{1}-\dots-n_{r}E_{r}$ be an ample line bundle on $X_{r}$ and let $\pi_{x}:\widetilde{X}_{r}(x)\to X_{r}$ denote the blow up of $X_{r}$ at $x$. We know that $\widetilde{C}^{2}<0$, where $\widetilde{C}$ the is strict transform of $C$. Let $t=\text{mult}_{x}C$. Since $\varepsilon(X_{r},L,x)<1$ we have $t>1$ and $L\cdot C\leq t-1$. By Lemma 4.1, $30>\widetilde{C}^{2}\geq-$ 3 so $C^{2}=t^{2}-i$, where $i=1$ or $2$ or $3$. Thus using the Hodge index theorem,
@@
So $(t-1)^{2}\geq t^{2}-i$, which reduces to $t\leq\frac{1+i}{2}$ for $i=1,2,3$. But this happens only if $t=1$ or $t=2$. As $\varepsilon(X_{r},L,x)<1$ we have $t=2$ and hence $\varepsilon(X_{r},L,x)=\frac{1}{2}$. ∎
@@
Consider $X_{r}$ as in Theorem 4.2 and suppose that $d\geq 1$, $t\geq 2$ and $m_{1},m_{2},\dots,m_{r}\geq 0$ be integers such that
(${d+2\choose 2}-\displaystyle\sum_{i=1}^{r}{m_{i}+1\choose 2}\leq{t\choose 2}.$
@@
Note that for $t\geq 2$, (${t+1\choose 2}-2\geq{t\choose 2}$. Hence Conjecture 4.4 follows from Conjecture 1.3.
@@
Let $X_{r}$ be as in Theorem 4.2. If Conjecture 4.4 is true then $\varepsilon(X_{r},L,x)\geq\frac{1}{2}$ for all ample line bundles $L$ on $X_{r}$ and for any $x\in X_{r}$.
@@
Let $\pi_{x},\widetilde{X}_{r}(x),E_{x}$ be as above and let $L=eH-n_{1}E_{1}-n-\dots-n_{r}E_{r}$. To show that $\varepsilon(X_{r},L,x)\geq\frac{1}{2}$, it is enough to show that $\pi_{x}^{\ast}L-\frac{1}{2}E_{x}$ is nef, i.e., $2\pi_{x}^{\ast}L-E_{x}$ is nef. Equivalently, for any reduced, irreducible curve $\widetilde{C}=dH-m_{1}E_{1}-\dots-m_{r}E_{r}-tE_{x}$ in $\widetilde{X}_{r}(x)$, we need to show that
$2(de-m_{1}n_{1}-\dots-m_{r}n_{r})-t\geq 0$
If $t=1$ clearly $2(de-m_{1}n_{1}-\dots-m_{r}n_{r})-t\geq 0$ If $\widetilde{C}=E_{i}\text{ or }\widetilde{C}=E_{i}-E_{x}$ clearly $(\pi_{x}^{\ast}L-\frac{1}{2}E_{x})\cdot\widetilde{C}\geq 0$.
Now assume $d\geq 1$ and $t\geq 2$. Since $\widetilde{C}$ is a reduced curve in $\widetilde{X}_{r}(x)$, $C=\pi_{x}(\widetilde{C})$ is a reduced curve in $X_{r}$. So using Conjecture 4.4 we conclude:
(${d+2\choose 2}-\displaystyle\sum_{i=1}^{r}{m_{i}+1\choose 2}>{t\choose 2}.$
Consider a very general point $y\in X_{r}$ and let $p_{r+1}:=\pi(y).$ Therefore, by dimension count, there is a curve $D$ in $\mathbb{P}^{2}$ such that $\text{mult}_{p_{i}}D\geq m_{i}$ for all $i=1,2,\dots,r$ and $\text{mult}_{p_{r+1}}D\geq t-1$.
@@
$2(de-m_{1}n_{1}-\dots-m_{r}n_{r})-t\geq 0$
This gives the required inequality
@@
One of the most important questions in the area of Seshadri constants concerns their irrationality, i.e., is there an ample line bundle on a smooth surface so that at some point in the surface, the Seshadri constant of that line bundle is irrational? This will be addressed in a later section. But now we propose another question:
Question 4.6.
@@
Clearly Theorem 4.2 implies that if Conjecture 1.3 is true, the above question has a negative answer for $\frac{p}{q}<\frac{1}{2}$. So it is enough to look at $\frac{p}{q}$ with $2p\geq q$. For this, it is enough to look for line bundles with Seshadri constant $\frac{p}{q}$, where $p$ is a prime number.
We give below a partial result in this direction.

### v3

@@ 4. Blow ups of $r\leq 8$ distinct points
[10, Proposition 2.6] Let $X_{r}$ be the blow up of $r\geq 1$ distinct points on $\mathbb{P}^{2}$ and let $L$ denote an ample line bundle on $X_{r}$. Assume $r\leq 8$, and for $x\in X_{r}$, let $\widetilde{X}_{r}(x)$ denote the blow up of $X_{r}$ at $x$. Then
@@
where $C$ is a curve on $X_{r}$ and $\Gamma$ is the set of all (1)(-1)-curves, (2)(-2)-curves, fixed components of $\big{|}-K_{\widetilde{X}_{r}(x)}\big{|}$ and reduced irreducible curves in the linear system $\big{|}-K_{\widetilde{X}_{r}(x)}\big{|}$ on $\widetilde{X}_{r}(x)$.
@@
If $C$ is a Seshadri curve on $X_{r}$ for an ample line bundle at a point $x\in X_{r}$, then $\widetilde{C}^{2}<0$, where $\widetilde{C}$ is the strict transform of $C$ on $\widetilde{X}_{r}(x)$. Note that by [15, Proposition 4.1 (ii)], for $r\leq 8$, $\widetilde{C}$ is either a (1)(-1)-curve or a (2)(-2)-curve or a fixed component of $\big{|}-K_{\widetilde{X}_{r}(x)}\big{|}$.
@@
Let $X$ be a smooth anti-canonical surface (i.e., $K_{X}$ is an effective divisor). If $(-K_{X})^{2}\geq 2$ we have,
$\varepsilon(X)\geq 1.$
@@
Let $L$ be an ample line bundle on $X_{r}$. Then as an application of the Hodge index theorem we have,
@@
So $K_{X}\cdot L\geq\sqrt{2}$. Also note that $K_{X}\cdot L$ is a positive integer. Thus we have $K_{X}\cdot L\geq 2$. Therefore by [12, Theorem III.1 (a)], $L$ is base point free. So $\varepsilon(X_{r},L,x)\geq 1$ for all $x\in X_{r}$. ∎
@@
For $r\leq 7$ we have,
$\varepsilon(X_{r})\geq 1.$
Next, we consider the $r=8$ case. First we prove the following lemma.
@@
Let $X_{8}$ be a blow up of $\mathbb{P}^{2}$ at 8 distinct points. If $L$ is an ample line bundle with $K_{X_{8}}\cdot L=1$ then $L=-K_{X_{8}}$.
@@
Let $L=eH-n_{1}E_{1}-\dots-n_{8}E_{8}$ be an ample line bundle with $K_{X_{8}}\cdot L=1$, i.e., $3e-\displaystyle\sum_{i=1}^{8}n_{i}=1$. If $L^{2}\geq 2$, the Hodge index theorem gives
@@
Let $X_{8}$ be a blow up of $\mathbb{P}^{2}$ at 8 distinct points $p_{1},p_{2},\dots,p_{8}$. Let $L$ be an ample line bundle on $X_{8}$. Suppose that $\varepsilon(X_{8},L,x)<1$ for some $x\in X_{8}$. Then the following hold:
@@
Suppose that $\varepsilon(X_{8},L,x)<1$ for an ample line bundle $L$ at some $x\in X_{8}$. By Proposition 4.1, the Seshadri constant is achieved by a curve $C$ on $X_{8}$ whose strict transform on $\widetilde{X}_{8}(x)$ is a (1)(-1)-curve or a (2)(-2)-curve or a fixed component of $|-K_{\widetilde{X}_{8}(x)}|$ or a curve in the linear system $|-K_{\widetilde{X}_{8}(x)}|$. Let $C=dH-m_{1}E_{1}-\dots-m_{8}E_{8}$ and $t=\text{mult}_{x}C>1$, i.e.,
@@
First, we show that the strict transform of $C$ on $\widetilde{X}_{8}(x)$ is not a (1)(-1)-curve or a (2)(-2)-curve. If $C$ is such a curve, we have $C^{2}=t^{2}-i$, where $i=1,2$. Since
@@
So $C$ is a singular curve whose strict transform is either a fixed component of $|-K_{\widetilde{X}_{8}(x)}|$ or a curve in the linear system $|-K_{\widetilde{X}_{8}(x)}|$. Since $C$ is a singular curve, we must have $d=3$. Hence $\widetilde{C}\in|-K_{\widetilde{X}_{8}(x)}|$. So $p_{1},\dots,p_{8}$ belong to the image of $C$ on $\mathbb{P}^{2}$. As $C$ is irreducible, it is clear that $\text{mult}_{x}C=2$. Since $\varepsilon(X_{8},L,x)<1$, the only possibility is
@@
A part of Theorem 4.6 has an alternate proof using Theorem 3.1. Note that there is a cubic passing through any distinct 8 points in $\mathbb{P}^{2}$. Theorem 3.1 gives that $\frac{1}{2}$ is a lower bound for the Seshadri constant of $X_{8}$. More precisely, using Lemma 4.5 we can see that $\varepsilon(X_{8},L)\geq 1$ for all ample line bundles $L\neq-K_{X_{8}}$ and $\varepsilon(X_{8},-K_{X_{8}})\geq\frac{1}{2}$ provided $K_{X_{8}}$ is an ample line bundle.
@@
By Theorem 4.6, the least Seshadri constant that can be achieved on $X_{8}$ is $\frac{1}{2}$. We now give an example in which $\frac{1}{2}$ is achieved at some point on a del Pezzo surface. Consider the smooth rational surface obtained by blowing up 8 smooth points $p_{1},p_{2},\dots p_{8}$ on a singular irreducible cubic on $\mathbb{P}^{2}$ such that no three lie on a line, no six lie on a conic, and if a singular cubic passes through all the points then $p_{i}$ are smooth points of that cubic for $1\leq i\leq 8$. Hence $K_{X_{8}}$ is ample by [4, Théorème 1]. Let $x$ denote the point in $X_{8}$ corresponding to the singular point on the cubic. Let $C$ denote the strict transform of the singular cubic containing the 8 points. From Theorem 4.6 it is clear that $\varepsilon(X_{8},-K_{X_{8}},x)\geq\frac{1}{2}$. Note that $\text{mult}_{x}C=2$ and $K_{X_{8}}\cdot C=1$. These give us:
@@
It is known that the number of (1)(-1) and (2)(-2)-curves on $\widetilde{X}_{r}(x)$ is finite for $r\leq 7$. Also for $r\leq 5$, we can see that curves whose strict transforms are (1)(-1) or (2)(-2)-curves on $\widetilde{X}_{r}(x)$ are smooth at $x$. The curves on $X_{r}$ whose strict transform is a fixed component of $|-K_{\widetilde{X}_{r}(x)}|$ and is not an element of $|-K_{\widetilde{X}_{r}(x)}|$ are also smooth. It is also clear that a curve $C$ on $X_{r}$ such that $\widetilde{C}$ is an element of $|-K_{\widetilde{X}_{r}(x)}|$ has non-negative self intersection if $r\leq 5$. Note the fact that if $C$ is a Seshadri curve, its strict transform $\widetilde{C}$ has negative self intersection on $\widetilde{X}_{r}(x)$. This shows that all possible Seshadri curves are smooth at $x$ when $r\leq 5$. Hence the Seshadri constant of every ample line bundles at every point of $X_{r}$ is an integer when $r\leq 5$. ∎
@@
It is shown in [2, Théorème 1.3] that there is a del Pezzo surface $X_{6}$ which is the blow up of six points on $\mathbb{P}^{2}$ such that no three lie on a line, no six lie on a conic, with fractional Seshadri constant. More precisely, we have $\varepsilon(X_{6},-K_{X_{6}},x)=\frac{3}{2}$ for some $x\in X_{6}$.
@@ 5. General blow ups of $\mathbb{P}^{2}$
We proceed with a similar setting as in the above section. Furthermore throughout this section we assume that $p_{1},p_{2},\dots,p_{r}$ are very general points in $\mathbb{P}^{2}$, where $r\geq 1$. In this section our goal is to find some lower bounds on Seshadri constants on $X_{r}$ assuming the Strong SHGH Conjecture (Conjecture 2.3) is true.
Following result classify all the possibilities of a self intersection negative curve on $\widetilde{X}_{r}(x)$ for any $x\in X_{r}$.
@@
For any point $x\in X_{r}$, let $\widetilde{X}_{r}(x)$ be the blow up of $X_{r}$ at x. Assume that Conjecture 2.3 is true. Let $C$ be a reduced, irreducible curve on $X_{r}$ and let $\widetilde{C}$ denote its strict transform on $\widetilde{X}_{r}(x)$. If $\widetilde{C}^{2}<0$ then $\widetilde{C}$ is one of the following holds:
@@
Let $C=dH-m_{1}E_{1}-\dots-m_{r}E_{r}$ and $t=\text{mult}_{x}C$. So $\widetilde{C}=dH-m_{1}E_{1}-\dots-m_{r}E_{r}-tE_{x}$, where $E_{x}$ is the exceptional divisor on $\widetilde{X}_{r}(x)$. If $d=0$, $\widetilde{C}=E_{i}$ or $\widetilde{C}=E_{i}-E_{x}$. Both are of the required type.
@@
${d+2\choose 2}-\displaystyle\sum_{i=1}^{r}{m_{i}+1\choose 2}>{t+1\choose 2}-2.$
@@
which is same as
$\frac{\widetilde{C}^{2}-K_{\widetilde{X}_{r}(x)}\cdot\widetilde{C}}{2}+1>-2.$ (5.1)
The Riemann-Roch theorem gives,
@@
Since $\widetilde{C}$ is an effective divisor on $\widetilde{X}_{r}(x)$, we get $h^{2}(\widetilde{X}_{r}(x),\widetilde{C})=0$. Also $\widetilde{C}^{2}<0$ implies $h^{0}(\widetilde{X}_{r}(x),\widetilde{C})=1$. Then (5.1) reduces to:
$h^{1}(\widetilde{X}_{r}(x),\widetilde{C})=\frac{\widetilde{C}^{2}-K_{\widetilde{X}_{r}(x)}\cdot\widetilde{C}}{2}>-3.$ (5.2)
@@
    From (5.2) we get, $h^{1}(\widetilde{X}_{r}(x),\widetilde{C})=\frac{K_{\widetilde{X}_{r}(x)}\cdot\widetilde{C}-\widetilde{C}^{2}}{2}=0$. i.e., $K_{\widetilde{X}_{r}(x)}\cdot\widetilde{C}=\widetilde{C}^{2}$.
    Now from adjunction formula we see that,
@@
    Similarly as in above cases, we conclude that types (3) or (5) or (6) can occur.
In fact Lemma 5.1 says that $C^{2}\geq-3$ for all reduced irreducible curve $C$ on $\widetilde{X}_{r}(x)$. The following is the main result of this section.
@@
Assume that Conjecture 2.3 is true. Let $\pi:X_{r}\to\mathbb{P}^{2}$ be the blow up of $\mathbb{P}^{2}$ at very general points $p_{1},p_{2},\dots,p_{r}$. If $\varepsilon(X_{r},L,x)<1$ for some ample line bundle $L$ and for some $x\in X_{r}$, then $\varepsilon(X_{r},L,x)=\frac{1}{2}$. In particular,
$\varepsilon(X_{r})\geq\frac{1}{2}.$
@@
Let $L$ be an ample line bundle on $X_{r}$ with $\varepsilon(X_{r},L,x)<1$ for some $x\in X_{r}$. Then $\varepsilon(X_{r},L,x)$ is achieved by an irreducible and reduced curve $C\subseteq X_{r}$, i.e.,
@@
We know that $\widetilde{C}^{2}<0$, where $\widetilde{C}$ is the strict transform of $C$. Let $t=\text{mult}_{x}C$. Then $\widetilde{C}^{2}=C^{2}-t^{2}$. Since $\varepsilon(X_{r},L,x)<1$ we have $t>1$ and $L\cdot C\leq t-1$. By Lemma 5.1 we have, $0>\widetilde{C}^{2}\geq-3$. Hence $C^{2}=t^{2}-i$, where $i=1$ or $2$ or $3$.
Now using the Hodge index theorem,
@@
So $(t-1)^{2}\geq t^{2}-i$, which reduces to $t\leq\frac{1+i}{2}$ for $i=1,2,3$. Hence $t=1$ or $t=2$. But $\varepsilon(X_{r},L,x)<1$ forces $t=2$ and hence $\varepsilon(X_{r},L,x)=\frac{1}{2}$. ∎
@@
Let $X_{r}$ be as in Theorem 5.2 and let $d\geq 1$, $t\geq 2$ and $m_{1},m_{2},\dots,m_{r}\geq 0$ be integers such that
${d+2\choose 2}-\displaystyle\sum_{i=1}^{r}{m_{i}+1\choose 2}\leq{t\choose 2}$.
@@
Note that for $t\geq 2$, ${t+1\choose 2}-2\geq{t\choose 2}$. Hence Conjecture 5.4 follows from Conjecture 2.3.
@@
Let $X_{r}$ be as in Theorem 5.2. If Conjecture 5.4 is true then $\varepsilon(X_{r})\geq\frac{1}{2}$.
@@
Let $\pi_{x},\widetilde{X}_{r}(x),E_{x}$ be as above. Let $L=eH-n_{1}E_{1}-n_{2}E_{2}-\dots-n_{r}E_{r}$ be an ample line bundle on $X_{r}$ and $x\in X_{r}$ be a point. To show that $\varepsilon(X_{r},L,x)\geq\frac{1}{2}$, it is enough to show that $\pi_{x}^{\ast}L-\frac{1}{2}E_{x}$ is nef, i.e., $2\pi_{x}^{\ast}L-E_{x}$ is nef. Equivalently, for any reduced, irreducible curve $\widetilde{C}=dH-m_{1}E_{1}-\dots-m_{r}E_{r}-tE_{x}$ on $\widetilde{X}_{r}(x)$, we need to show that
$2(de-m_{1}n_{1}-\dots-m_{r}n_{r})-t\geq 0.$
Since $L$ is ample, we have $L\cdot C=de-m_{1}n_{1}-\dots-m_{r}n_{r}\geq 1$. Hence if $t=1$ clearly $2(de-m_{1}n_{1}-\dots-m_{r}n_{r})-t\geq 0.$ Also if $d=0$ we have, $\widetilde{C}=E_{i}\text{ or }\widetilde{C}=E_{i}-E_{x}$ clearly $(\pi_{x}^{\ast}L-\frac{1}{2}E_{x})\cdot\widetilde{C}\geq 0$.
[…]
${d+2\choose 2}-\displaystyle\sum_{i=1}^{r}{m_{i}+1\choose 2}>{t\choose 2}$.
Consider a very general point $y\in X_{r}$ and let $p_{r+1}:=\pi(y).$ Therefore, by dimension count, there is a curve $D$ on $\mathbb{P}^{2}$ such that $\text{mult}_{p_{i}}D\geq m_{i}$ for all $i=1,2,\dots,r$ and $\text{mult}_{p_{r+1}}D\geq t-1$.
@@
$2(de-m_{1}n_{1}-\dots-m_{r}n_{r})-t\geq 0.$
This gives the required inequality:
@@
One of the most important questions in the area of Seshadri constants concerns their irrationality, i.e., is there an ample line bundle on a smooth surface so that at some point in the surface, the Seshadri constant of that line bundle is irrational? We propose the following question:
Question 5.6.
@@
Clearly Theorem 5.2 implies that if Conjecture 2.3 is true, Question 5.6 has a negative answer for $\frac{p}{q}<1$ unless $\frac{p}{q}=\frac{1}{2}$. It is enough to look at $\frac{p}{q}$ with $p\geq q$. For this, it is enough to look for line bundles with Seshadri constant $\frac{p}{q}$, where $p$ is a prime number as $\varepsilon(X,nL,x)=n\varepsilon(X,L,x).$
We give below a partial result in this direction. The following result shows all integers are attained as Seshadri constants.